let f(x)=x^4-2x^2. Find the all c (where c is the interception on the x-axis) in the interval (-2,2) such that f'(x)=0 using Rolle's theorem.
1
Expert's answer
2019-05-14T10:22:00-0400

Solution. Using Rolle's theorem. Suppose that a function f(x)  is continuous on the closed interval [a;b] and differentiable on the open interval (a;b). Then if f(a)=f(b), then there exists at least one point c in the open interval (a;b) for which f'(c)=0.

Function  f(x)=x^4-2x^2  is continuous on the closed interval [-2;2] and differentiable on the open interval (-2;2).


"f(-2)=(-2)^4-2*(-2)^2=16-8=8"

"f(2)=2^4-2*2^2=16-8=8""f(-2)=f(2)"

On the other hand


"f'(x)=4x^3-4x."

Hence get equation


"4x^3-4x=0"

Solve the equation


"4x(x^2-1)=0"

Roots of the equation


"x_1=0 \isin (-2;2)"

"x_2=-1 \isin (-2;2)"

"x_3=1 \isin (-2;2)"

Answer. -1, 0, 1.